If sin x = 4/5 and x is in the second quadrant, find the remaining 5 trig values for the angle x.
Answer by lwsshak3(11628)

If sin x = 4/5 and x is in the second quadrant, find the remaining 5 trig values for the angle x.
***
sinx=4/5 (working with a (3-4-5) reference right triangle in quadrant II in which sin>0, cos<0
cosx=-3/5
cscx=5/4
secx=-5/3
tanx=-4/3
cotx=-3/4